OpenStudy (anonymous):

The U.S. Senate has 100 members. Suppose that it is desired to place each senator on exactly 1 of 7 possible committees. The first committee has 22 members, the second has 13, the third has 10, the fourth has 5, the fifth has 16, and the sixth and seventh have 17 apiece. In how many ways can these committees be formed?

OpenStudy (anonymous):

Ohh I was typing a long equation and I lost it.

OpenStudy (anonymous):

I got this:

OpenStudy (anonymous):

\[C(100,22)\cdot C(78,13)\cdot C(65,10)\cdot C(55,5)\cdot\]

OpenStudy (anonymous):

\[C(50,16)\cdot C(34,17)\cdot C(17,17)\]

OpenStudy (anonymous):

\[=1.156\times10^{76}\]

OpenStudy (anonymous):

I think Its wrong because it is a very big number.

OpenStudy (anonymous):

What do you think?

OpenStudy (across):

If with \(C(x,y)\) you mean \(x\) choose \(y\), then your thought process seems correct.\[\]

OpenStudy (across):

Or, more formally,\[C(x,y)=\frac{x!}{y!(x-y)!}.\]

OpenStudy (anonymous):

So that means my result is correct. I'm very surprised then. Thank you across.